# On the application of Lorentz-Finsler geometry to model wave propagation

### v1

@@ 2.5.1 Time-independent case
Assume first that the speed profile of the wave is time-independent, i.e. $V=V(x,v)$, so that it provides a proper Finsler metric $F_{x}(v)=\frac{||v||_{h_{x}}}{V(x,v)}$. In this case, note that we can directly integrate (3) to obtain the total traveltime spent by the wave ray when following $\sigma$:
@@
which coincides with the $F$-length of $\sigma$, defined in (6), i.e. the (time-independent) Finsler metric $F$ acts in a similar way as $h$ in § 2.1, but instead of measuring distances, it directly measures traveltimes. According to Fermat’s principle, $\sigma$ must be a critical point of $T$ among all the curves in $\mathcal{N}_{x,y}$, which means that wave rays follow the pregeodesic curves of $F$.
@@ 2.5.2 Time-dependent case
When the speed profile is time-dependent, the differential equation (3) is not so easy to solve for the time function $t(s)$, since now it appears inside $V$ [^10] Although we will not follow this approach here, calculus of variations can be directly applied to (3), leading to a generalized version of the Euler-Lagrange equations for $\sigma$ to be a critical point of the traveltime (see, e.g. [12, § 2] and [20, § 3]). Moreover, as previously discussed, $F$ is no longer a standard metric whose pregeodesics can be defined in the usual way. Instead, we can consider the more general spacetime viewpoint constructed in § 2.4, using the Lorentz-Finsler metric $G$ on $M=\mathds{R}\times N$. Note that the curve $\gamma(s)=(t(s),\sigma(s))$ in the spacetime directly incorporates the measurement of the traveltime in its first component. Also, we have already pointed out that (3) is equivalent to (9), i.e. $\gamma$ must be lightlike when representing a wave path. This way, traveling between two points $x,y\in N$ in the space is equivalent for the wave ray to go from $(0,x)\in M$ (choosing $t=0$ as the initial time) to the vertical line $\ell_{y}:t\mapsto(t,y)$ following a lightlike curve in the spacetime.
Therefore, if we consider the set $\mathcal{N}_{x,\ell_{y}}$ of all the (piecewise smooth) lightlike curves $\gamma:[a,b]\rightarrow M$ from $\gamma(a)=(0,x)$ to $\gamma(b)\in\ell_{y}$, the traveltime functional is given by
@@
and Fermat’s principle tells us that a wave path in the spacetime must be a critical point of $T$ among all the curves in $\mathcal{N}_{x,\ell_{y}}$. It turns out that these critical points are precisely the lightlike pregeodesics of $G$ (see [23, Theorem 4.2]).
@@ 2.6 Initial wavefront
From the previous study, we have come to the conclusion that the path taken by a wave ray between two fixed points is a lightlike pregeodesic of $G$. Let us assume now, according to the initial formulation of the problem, that the wave departs not from a single point, but from an entire subset $S\subset N$ acting as an initial wavefront at time $t=0$. We will assume that this initial wavefront is a compact (smooth embedded) hypersurface of $N$ (if $n=2$, then $S$ is a simple closed curve). [^11] This can be generalized to the case when the initial wavefront $S$ is a submanifold of $N$ of arbitrary codimension (see [14]). The wave paths then have to be critical points of the traveltime functional (13), but now in the broader set $\mathcal{N}_{S,\ell_{y}}$ of (piecewise smooth) lightlike curves $\gamma:[a,b]\rightarrow$ from $\sigma(a)\in S$ to $\gamma(b)\in\ell_{y}$. The additional requirement for $\gamma$ to be critical is that it must depart orthogonally from $S$ (see [14, § 4.1]).
@@
Given any point $x\in S$, there are exactly two $F$-orthogonal directions to $S$ (or, equivalently, two $G$-orthogonal lightlike directions to $S$; see [1, Proposition 5.2]): one points to the interior of $S$, and the other to the exterior. Here we will focus on the outward trajectories—which are usually the most interesting ones from a practical viewpoint—, although the results we will present directly apply to the inward trajectories simply by replacing “outward” with “inward”.
@@ 2.7 Geodesic equations
$\begin{split}G_{p}(\hat{v})&=G(x^{0},\ldots,x^{n},v^{0},\ldots,v^{n}),\\ \gamma(s)&=(\gamma^{0}(s),\ldots,\gamma^{n}(s)),\\ \end{split}$
@@
Also, Einstein’s summation convention will be used, summing the indices that appear up and down. Indices $i,j,k,r$ run from 00 to $n$.
@@
where $\hat{g}_{ij}(p,\hat{v})$ refers to the coefficients of the coordinate matrix of the fundamental tensor $g^{G_{p}}_{\hat{v}}$, given by (11). Taking the derivative with respect to the parameter $s$ in the first equation, and omitting the explicit reference to $s$ to avoid cluttering:
@@
where $\hat{\gamma}_{\ ij}^{k}(p,\hat{v})$ are the formal Christoffel symbols of $G$, given by
$\hat{\gamma}_{\ ij}^{k}(p,\hat{v})\coloneqq\frac{1}{2}g^{kr}(p,\hat{v})\left(\frac{\partial\hat{g}_{rj}}{\partial x^{i}}(p,\hat{v})+\frac{\partial\hat{g}_{ri}}{\partial x^{j}}(p,\hat{v})-\frac{\partial\hat{g}_{ij}}{\partial x^{r}}(p,\hat{v})\right),$ (16)
@@
Observe that (15) is an ordinary differential equation (ODE) system and thus, geodesics are univocally determined by their initial point and velocity. So, fixing $x\in S$ and $v\bot_{F}S$, there exists a unique lightlike geodesic $\gamma:[a,b]\rightarrow M$ (assuming $b$ is maximal), written as $\gamma(s)=(t(s),\sigma(s))$, such that $\gamma(a)=(0,x)$ and $\dot{\sigma}(a)=v$ (recall Remark 2). This curve $\gamma$ is then a critical point of the traveltime functional (13) on $\mathcal{N}_{S,\ell_{\sigma(s)}}$, for any $s\in(a,b]$.
@@ 3 The model
We now have all the ingredients we need to answer affirmatively to the initial question and compute the entire wave propagation. To this end, we will first construct the theoretical model that provides the evolution of the wavefront over time, and then we will see how to implement it in practice.
@@ 3.1 The wave trajectories
In the previous section, we have identified the paths taken by individual wave rays as the lightlike pregeodesics of a Lorentz-Finsler metric $G$, defined by (8), departing $F$-orthogonally to the initial wavefront $S$. Although we have been working with arbitrary parametrizations so far—in order to show explicitly the importance of the path itself, rather than the specific parametrization we choose—in practice it will be more convenient and natural to parametrize the curves by the time coordinate $t$. This leads to the following definition.
@@
Any $t$-parametrized curve $\sigma:[0,\tau]\rightarrow N$ such that $\gamma(t)=(t,\sigma(t))$ is a lightlike pregeodesic of $G$ with $\sigma(0)\in S$ and $\dot{\sigma}(0)\bot_{F}S$ will be called a wave trajectory.
@@
which is equivalent to $\dot{\sigma}(t)\in\Sigma_{t,\sigma(t)}$. From Remarks 3 and 4, note that a wave trajectory is uniquely determined by its initial position on the initial wavefront, so long we only consider outward trajectories. The idea then is to encompass all these trajectories in a single map.
@@ 3.2 The wavefront
Until now we have studied individual wave trajectories, but not how they generate the final wavefront. The way the wave behaves to provide this wavefront is governed by Huygens’ envelope principle (see, e.g. [21, § 2] and [14, § 3.3]):
The wavefront $\mathcal{W}_{\tau}$ at time $t=\tau$ is given by the envelope of all the possible wave trajectories departing from a previous wavefront $\mathcal{W}_{t_{0}}$ at time $t=t_{0}$ and running an interval $\tau-t_{0}$.
In the end, the trajectories that remain in the wavefront (i.e. those whose endpoint is an actual point in the wavefront) are the ones that, not only are critical points of the traveltime functional, but in fact global minima (see [14, § 4.1]). Indeed, if a wave trajectory $\sigma:[t_{0},\tau]\rightarrow N$ does not minimize the traveltime from the wavefront $\mathcal{W}_{t_{0}}$ to the endpoint $\sigma(\tau)$, it means that there exists another wave trajectory from $\mathcal{W}_{t_{0}}$ that reaches $\sigma(\tau)$ in a shorter time. Therefore, $\sigma(\tau)$ does not belong to the envelope of all the possible trajectories departing from $\mathcal{W}_{t_{0}}$, which means that it cannot be a point in $\mathcal{W}_{\tau}$. This motivates the following definition.
@@
We say that a wave trajectory $t\mapsto f(t,x_{0})$ is time-minimizing at time $t=\tau>0$ if for any fixed $t_{0}\in(0,\tau]$, any other wave trajectory $t\mapsto f(t,x_{1})$ such that $f(t_{0},x_{0})=f(t_{1},x_{1})$ satisfies that $t_{0}\leq t_{1}$. [^12] Essentially, this means that $t\mapsto(t,f(t,x_{0}))$ is a global minimum of the traveltime functional (13) on $\mathcal{N}_{S,\ell_{f(t_{0},x_{0})}}$, for any $t_{0}\in(0,\tau]$.
@@
and the following key result ensures that every wave trajectory remains in the wavefront, at least for a small time (see [14, Theorem 4.8]). [^13] In fact, the evolution of the wavefront over (a small) time generates a smooth lightlike hypersurface in the spacetime, as described in [13].
@@
This means that the wavefront $\mathcal{W}_{\tau}$ is directly given by $S\ni x\mapsto f(\tau,x)$, for any $\tau\leq\varepsilon$. Beyond this time, however, there might be trajectories in the wavemap that no longer provide points in the wavefront. A point $x\in N$ beyond which a wave trajectory is no longer time-minimizing is called a cut point (see [14, § 4.1] and [15, § 4.1]). These points are usually very interesting from a practical perspective, not only because they allow us to accurately describe the wavefront, but also because they mark regions where several wave trajectories tend to converge—e.g. in the case of wildfires, these regions can become extremely dangerous for firefighters, so it is important to detect them beforehand (see the comments in [15, § 5.2]).
@@ 3.3 The implementation
Lastly, we will see how this model can be implemented in practice. Assuming we know the speed profile $V(t,x,v)$—and thus, the (time-dependent) Finsler metric $F$ given by (4)—and the initial wavefront $S$, one has to implement the following steps:
@@ 3.3.1 Step 1: The initial wavefront
so that the vectors $\{\frac{\partial\alpha}{\partial\theta^{1}}(\theta),\ldots,\frac{\partial\alpha}{\partial\theta^{n-1}}(\theta)\}$ form a basis of $T_{\alpha(\theta)}S$, for all $\theta\in J$. This parametrization can always be obtained, at least locally—simply taking local coordinates on $S$.
@@ 3.3.2 Step 2: The orthogonality equations
where the fundamental tensor $g_{v}^{F_{t,x}}$ of $F$ is given by (5). This is, of course, not enough to determine the $n$ components $v_{l}=(v_{l}^{1},\ldots,v_{l}^{n})$. From Remark 3, we know that there are always two $F$-orthogonal directions to $S$ and we are looking for the one pointing outwards. Therefore, we need to fix the length of $v_{l}$ and its orientation. Regarding the length, (17) tells us that $v_{l}$ must be $F$-unitary:
@@
Regarding the orientation, we need to ensure that the determinant
@@ 3.3.3 Step 3: The pregeodesic equations
For each fixed $l\in\{1,\ldots,m\}$, the wave trajectory $t\mapsto f(t,\theta_{l})$ has to be the unique lightlike pregeodesic of $G$ with initial velocity $v_{l}$ (recall Definitions 3 and 4). We have already derived in § 2.7 the geodesics equations (15) to obtain the unique lightlike geodesic $\tilde{\gamma}(s)=(t(s),\tilde{\sigma}(s))$ departing from a fixed initial point and velocity. If we reparametrize $\tilde{\gamma}$ by the time coordinate, so that
@@
it is easy to check that the geodesic equations transform into the following $t$-parametrized pregeodesic equations (here indices $i,j,k,r$ run from 00 to $n$ and we are omitting the explicit reference to the parameter $t$ of $\gamma$); see [14, § 4.2]:
@@
However, notice that $\gamma^{0}(t)=t$ is already determined, so from the definition of $\hat{\gamma}_{\ ij}^{k}(p,\hat{v})$ in (16) and the relationship between the fundamental tensors of $F$ and $G$ in (11), we can write (21) in terms of the spatial curve $\sigma(t)$ and the (time-dependent) Finsler metric $F$ (here indices $\mu,\nu,\xi,\zeta$ run from $1$ to $n$ and we omit again the explicit reference to the parameter $t$ of $\sigma$):
$\ddot{\sigma}^{\xi}+g^{\xi\mu}(t,\sigma,\dot{\sigma})\frac{\partial g_{\mu\nu}}{\partial t}(t,\sigma,\dot{\sigma})\dot{\sigma}^{\nu}+\dot{\sigma}^{\mu}\dot{\sigma}^{\nu}\left(\gamma^{\xi}_{\ \mu\nu}(t,\sigma,\dot{\sigma})+\frac{1}{2}\frac{\partial g_{\mu\nu}}{\partial t}(t,\sigma,\dot{\sigma})\dot{\sigma}^{\xi}\right)=0.$ (22)
@@
$\gamma^{\xi}_{\ \mu\nu}(t,x,y)\coloneqq\frac{1}{2}g^{\xi\zeta}(t,x,y)\left(\frac{\partial g_{\zeta\nu}}{\partial x^{\mu}}(t,x,y)+\frac{\partial g_{\zeta\mu}}{\partial x^{\nu}}(t,x,y)-\frac{\partial g_{\mu\nu}}{\partial x^{\zeta}}(t,x,y)\right),$
@@
In conclusion, in this step we need to solve the ODE system (22) for $\sigma(t)$ in an interval $t\in[0,\tau]$, with the initial conditions $\sigma(0)=\alpha(\theta_{l})$ and $\dot{\sigma}(0)=v_{l}$. This is the wave trajectory $t\mapsto f(t,\theta_{l})=\sigma(t)$. Repeating steps 2 and 3 for every initial point $\{\theta_{l}\}_{l=1}^{m}$, we obtain the complete discrete wavemap for $t\in[0,\tau]$. If computing the individual wave trajectories is enough for our purposes, we can stop here. If, on the other hand, an explicit representation of the wavefront is needed, we have to proceed to the next step.

### v2

@@ 2.5.1 Time-independent case
[…]
which coincides with the $F$-length of $\sigma$, defined in (6), i.e. the (time-independent) Finsler metric $F$ acts in a similar way as $h$ in § 2.1, but instead of measuring distances, it directly measures traveltimes. According to Fermat’s principle, $\sigma$ must be a critical point of $T$ among all curves in $\mathcal{N}_{x,y}$, which means that wave rays follow the pregeodesic curves of $F$.
@@ 2.5.2 Time-dependent case
When the speed profile is time-dependent, the differential equation (3) is not so easy to solve for the time function $t(s)$, since now it appears inside VV—observe that we cannot integrate (3) directly unless we know $t(s)$ in advance. [^10] Although we will not follow this approach here, calculus of variations can be directly applied to (3), leading to a generalized version of the Euler-Lagrange equations for $\sigma$ to be a critical point of the traveltime (see e.g. [12, § 2] and [20, § 3]). Moreover, as previously discussed, $F$ is no longer a standard metric whose pregeodesics can be defined in the usual way. Instead, we can consider the more general spacetime viewpoint constructed in § 2.4, using the Lorentz-Finsler metric $G$ on $M=\mathds{R}\times N$. Note that the curve $\gamma(s)=(t(s),\sigma(s))$ in the spacetime directly incorporates the measurement of the traveltime in its first component. Also, we have already pointed out that (3) is equivalent to (9), i.e. $\gamma$ must be lightlike when representing a wave path. This way, traveling between two points $x,y\in N$ in the space is equivalent for the wave ray to go from $(0,x)\in M$ (choosing $t=0$ as the initial time) to the vertical line $\ell_{y}:t\mapsto(t,y)$ following a lightlike curve in the spacetime.
Therefore, if we consider the set $\mathcal{N}_{x,\ell_{y}}$ of all (piecewise smooth) lightlike curves $\gamma:[a,b]\rightarrow M$ from $\gamma(a)=(0,x)$ to $\gamma(b)\in\ell_{y}$, the traveltime functional is given by
@@
and Fermat’s principle tells us that a wave path in the spacetime must be a critical point of $T$ among all curves in $\mathcal{N}_{x,\ell_{y}}$. It turns out that these critical points are precisely the lightlike pregeodesics of $G$ (see [23, Theorem 4.2]).
@@ 2.6 Initial wavefront
From the previous study, we have come to the conclusion that the path taken by a wave ray between two fixed points is a lightlike pregeodesic of $G$. Let us assume now, according to the initial formulation of the problem, that the wave departs not from a single point, but from an entire subset $S\subset N$ that acts as the initial wavefront at time $t=0$. We will assume that this initial wavefront is a compact (smooth embedded) hypersurface of $N$ (if $n=2$, then $S$ is a simple closed curve). [^11] This can be generalized to the case when the initial wavefront $S$ is a submanifold of $N$ of arbitrary codimension (see [14]). The wave paths then have to be critical points of the traveltime functional (12), but now in the broader set $\mathcal{N}_{S,\ell_{y}}$ of (piecewise smooth) lightlike curves $\gamma:[a,b]\rightarrow M$ from $\sigma(a)\in S$ to $\gamma(b)\in\ell_{y}$. This imposes an additional requirement for $\gamma$ to be critical: $\gamma$ must depart orthogonally from $S$ (see [14, § 4.1]).
@@
Given any point $x\in S$, there are exactly two $F$-orthogonal directions to $S$ (or, equivalently, two $G$-orthogonal lightlike directions to $S$; see [1, Proposition 5.2]): one points to the interior of $S$, and the other to the exterior. Here we will focus on the outward trajectories—which are usually the most interesting ones from a practical viewpoint—although the results we will present directly apply to the inward trajectories simply by replacing “outward” with “inward”.
@@ 2.7 Geodesic equations
[…]
@@
Also, Einstein’s summation convention will be used, summing the indices that appear up and down. Indices $i,j,k,r$ run from 0 to $n$.
@@
[…]
@@
[…]
@@
Observe that (14) is an ordinary differential equation (ODE) system and thus, geodesics are univocally determined by their initial position and velocity. So, fixing $x\in S$ and $v\bot_{F}S$, there exists a unique lightlike geodesic $\gamma:[a,b]\rightarrow M$ (assuming $b$ is maximal), written as $\gamma(s)=(t(s),\sigma(s))$, such that $\gamma(a)=(0,x)$ and $\dot{\sigma}(a)=v$ (recall Remark 2). This curve $\gamma$ is then a critical point of the traveltime functional (12) on $\mathcal{N}_{S,\ell_{\sigma(s)}}$, for any $s\in(a,b]$.
@@ 3 The model
We now have all the ingredients needed to answer affirmatively the initial question posed at the beginning of § 2 and compute the entire wave propagation. To this end, we will first construct the theoretical model that provides the evolution of the wavefront over time, and then we will outline how to implement it in practice.
@@ 3.1 The wave trajectories
In the previous section, we have identified the paths taken by individual wave rays as lightlike pregeodesics of a Lorentz-Finsler metric $G$, defined by (8), departing $F$-orthogonally from the initial wavefront $S$. Although we have been working with arbitrary parametrizations so far—in order to show explicitly the importance of the path itself, rather than the specific parametrization we choose—in practice it will be more convenient and natural to parametrize the curves by the time coordinate $t$. This leads to the following definition.
@@
Any $t$-parametrized curve $\sigma:[0,\tau]\rightarrow N$ such that $\gamma(t)=(t,\sigma(t))$ is a lightlike pregeodesic of $G$, with $\sigma(0)\in S$ and $\dot{\sigma}(0)\bot_{F}S$, will be called a wave trajectory.
@@
which is equivalent to $\dot{\sigma}(t)\in\Sigma_{t,\sigma(t)}$. From Remarks 3 and 4, note that a wave trajectory is uniquely determined by its initial position on the initial wavefront, provided that we only consider outward trajectories. The idea then is to encompass all these trajectories in a single map.
@@ 3.2 The wavefront
Until now we have studied individual wave trajectories, but not how they generate the final wavefront. The way the wave behaves to provide this wavefront is governed by Huygens’ envelope principle (see e.g. [21, § 2] and [14, § 3.3]):
The wavefront $\mathcal{W}_{\tau}$ at time $t=\tau$ is given by the envelope of all possible wave trajectories departing from a previous wavefront $\mathcal{W}_{t_{0}}$ at time $t=t_{0}$ and running an interval $\tau-t_{0}$.
In the end, trajectories that remain in the wavefront (i.e., those whose endpoint is an actual point in the wavefront) are the ones that not only are critical points of the traveltime functional but in fact global minima (see [14, § 4.1]). Indeed, if a wave trajectory $\sigma:[t_{0},\tau]\rightarrow N$ does not minimize the traveltime from the wavefront $\mathcal{W}_{t_{0}}$ to the endpoint $\sigma(\tau)$, it means that there exists another wave trajectory from $\mathcal{W}_{t_{0}}$ that reaches $\sigma(\tau)$ in a shorter time. Therefore, $\sigma(\tau)$ does not belong to the envelope of all possible trajectories departing from $\mathcal{W}_{t_{0}}$, which means that it cannot be a point in $\mathcal{W}_{\tau}$. This motivates the following definition.
@@
We say that a wave trajectory $t\mapsto f(t,x_{0})$ is time-minimizing at time $t=\tau>0$ if, for any fixed $t_{0}\in(0,\tau]$, any other wave trajectory $t\mapsto f(t,x_{1})$ such that $f(t_{0},x_{0})=f(t_{1},x_{1})$ satisfies that $t_{0}\leq t_{1}$. [^12] Essentially, this means that $t\mapsto(t,f(t,x_{0}))$ is a global minimum of the traveltime functional (12) on $\mathcal{N}_{S,\ell_{f(t_{0},x_{0})}}$, for any $t_{0}\in(0,\tau]$.
@@
and the following key result ensures that every wave trajectory remains in the wavefront, at least for a small time (see [14, Theorem 4.8]). [^13] In fact, the evolution of the wavefront over (a small) time generates a smooth lightlike hypersurface of the spacetime, as described in [13].
@@
This means that the wavefront $\mathcal{W}_{\tau}$ is directly given by $S\ni x\mapsto f(\tau,x)$, for any $\tau\leq\varepsilon$. Beyond this time, however, there might be trajectories in the wavemap that no longer provide points in the wavefront. A point $x\in N$ beyond which a wave trajectory is no longer time-minimizing is called a cut point (see [14, § 4.1] and [15, § 4.1]). These points are usually very interesting from a practical perspective, not only because they allow us to accurately describe the wavefront but also because they mark regions where several wave trajectories tend to converge—e.g., in the case of wildfires, these regions can become extremely dangerous for firefighters, so it is important to detect them beforehand (see the comments in [15, § 5.2]).
@@ 3.3 The implementation
Lastly, we will see how this model can be implemented in practice. Assuming we know the speed profile V(t,x,v)V(t,x,v)—and thus, the (time-dependent) Finsler metric $F$ given by (4)—and the initial wavefront $S$, one has to implement the following steps:
@@ 3.3.1 Step 1: The initial wavefront
so that the vectors $\{\frac{\partial\alpha}{\partial\theta^{1}}(\theta),\ldots,\frac{\partial\alpha}{\partial\theta^{n-1}}(\theta)\}$ form a basis of $T_{\alpha(\theta)}S$, for all $\theta\in J$. This parametrization can always be obtained, at least locally—simply by taking local coordinates on $S$.
@@ 3.3.2 Step 2: The orthogonality equations
where the fundamental tensor $g_{v}^{F_{t,x}}$ of $F$ is given by (5). This is, of course, not enough to determine the $n$ components of $v_{l}=(v_{l}^{1},\ldots,v_{l}^{n})$. We will also need to fix the length of $v_{l}$ and its orientation. Regarding the length, (16) tells us that $v_{l}$ must be $F$-unitary:
@@
Regarding the orientation, from Remark 3 we know that there are always two $F$-orthogonal directions to $S$ and we are looking for the one pointing outwards. Therefore, we need to ensure that the determinant
@@ 3.3.3 Step 3: The pregeodesic equations
For each fixed $l\in\{1,\ldots,m\}$, the wave trajectory $t\mapsto f(t,\theta_{l})$ has to be the unique lightlike pregeodesic of $G$ with initial velocity $v_{l}$ (recall Definitions 3 and 4). We have already derived in § 2.7 the geodesics equations (14) to obtain the unique lightlike geodesic $\tilde{\gamma}(s)=(t(s),\tilde{\sigma}(s))$ departing from a fixed initial position and velocity. If we reparametrize $\tilde{\gamma}$ by the time coordinate, so that
@@
it is easy to check that the geodesic equations transform into the following $t$-parametrized pregeodesic equations (here indices $i,j,k,r$ run from 0 to $n$ and we are omitting the explicit reference to the parameter $t$ of $\gamma$); see [14, § 4.2]:
@@
[…]
$\ddot{\sigma}^{\xi}+g^{\xi\mu}(t,\sigma,\dot{\sigma})\frac{\partial g_{\mu\nu}}{\partial t}(t,\sigma,\dot{\sigma})\dot{\sigma}^{\nu}+\dot{\sigma}^{\mu}\dot{\sigma}^{\nu}\left(\gamma^{\xi}_{\ \mu\nu}(t,\sigma,\dot{\sigma})+\frac{1}{2}\frac{\partial g_{\mu\nu}}{\partial t}(t,\sigma,\dot{\sigma})\dot{\sigma}^{\xi}\right)=0,$ (21)
@@
$\gamma^{\xi}_{\ \mu\nu}(t,x,v)\coloneqq\frac{1}{2}g^{\xi\zeta}(t,x,v)\left(\frac{\partial g_{\zeta\nu}}{\partial x^{\mu}}(t,x,v)+\frac{\partial g_{\zeta\mu}}{\partial x^{\nu}}(t,x,v)-\frac{\partial g_{\mu\nu}}{\partial x^{\zeta}}(t,x,v)\right),$
@@
In conclusion, in this step we need to solve the ODE system (21) for $\sigma(t)$ in an interval $t\in[0,\tau]$, with the initial conditions $\sigma(0)=\alpha(\theta_{l})$ and $\dot{\sigma}(0)=v_{l}$. This provides the wave trajectory $t\mapsto f(t,\theta_{l})=\sigma(t)$. Repeating steps 2 and 3 for every initial point $\{\theta_{l}\}_{l=1}^{m}$, we obtain the complete discrete wavemap for $t\in[0,\tau]$. If computing the individual wave trajectories is enough for our purposes, we can stop here. If, on the other hand, an explicit representation of the wavefront is needed, we have to proceed to the next step.